# Reassessing quantum-thermodynamic enhancements in continuous thermal machines

### v2

@@ VI.3 Noise-induced coherent machine
Notice that in the case of symmetric rates for the cold reservoir, $\gamma_{\mathrm{c}}^{\mathrm{a}}=\gamma_{\mathrm{c}}^{\mathrm{b}}$, this expression becomes the one presented in Ref. [69], but differs otherwise. By introducing this change, we successfully decouple the state |$\ket{\beta}$ from the state |$\ket{0}$, resulting in a machine with only one explicit collective transition, as illustrated in Fig. 5 (see Appendix for more detailed information).
Following Sec. IV.2, the classical equivalent of the NIC machine includes an extra stochastic transition between the degenerate levels reading:
together with corrections to the rates of the collective transitions
for $k=\alpha,\beta$. By examining the above equations, we find that the corrections to the spontaneous emission rates can make these rates eventually negative, which would correspond to a nonphysical situation. Consequently, the construction of the classical equivalent for noise-induced coherence is limited to scenarios that ensure positive rates in Eq. (39). For the NIC machine analyzed here, this happens when
We obtain analytically the fluctuation ratio $\mathcal{R}$ in Eq. (31) for the NIC machine and its classical equivalent, taking again $J_{\mathrm{out}}=\dot{W}$. By evaluating the expression for $10^{6}$ different parameter configurations within the region where the classical equivalent machine can be defined, we show in Fig. 6 the distribution of $\mathcal{R}$ together with the corresponding distribution of the TUR ratio $\mathcal{Q}_{\mathrm{he}}$ values (see inset). Our results show that $\mathcal{R}\geq 0$ for some configurations, pointing to quantum-thermodynamic advantages in a similar way than the two previous examples. However, here we also get configurations leading to $\mathcal{R}<0$. For that configurations the quantum-coherent machine becomes more noisy than its classical equivalent, hence leading to a detrimental role for the coherence. Indeed for some configurations the NIC machine reaches noise levels up to $1\%$ higher than its classical counterpart. Moreover, by looking at the inset we observe that, as expected, the TUR is not violated for all choices of parameters ($\mathcal{Q}_{\mathrm{he}}\geq 2$ always) in accordance to previous studies about TUR violations in similar models [68].
Figure 6: Histogram of sampled values of $\mathcal{Q}$ (inset plots) and the ratio between the fluctuations of the system and its classical equivalent $R$ (outset plots). The values corresponds to an exploration of the following region in the parameters space of the system: $\beta_{\mathrm{w}}\rightarrow 0$, $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{h}}/\beta_{\mathrm{c}}\in[0,1]$, $\epsilon_{2}=5$, $\epsilon_{1}\in[0.1,4.9]$, $\gamma_{\mathrm{h}}^{(\mathrm{a}/\mathrm{b})}\in[10^{-5},10^{-2}]$ and $\gamma_{\mathrm{c}}^{(\mathrm{a}/\mathrm{b})}=10^{-3}$ (energies are given in $k_{\rm B}T_{\rm c}=1$ units).
Figure 7: Violations of the TUR and two dimensional multi-objective optimization for power $\langle\dot{W}\rangle$, power fluctuations ${\rm Var}[\dot{W}]$ and efficiency $\eta$ for the Coherent three-level amplifier. Points are generated from $10^{5}$ different configurations in the parameter space $\omega_{\mathrm{d}}\in[0.1,4.9]$, $\gamma_{\mathrm{h}/\mathrm{c}}\in[10^{-5},10^{-2}]$ and $g\in[10^{-5},10^{-3}]$ and fixed $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{h}}=0.1$, $\epsilon_{2}=5$ (energies are given in $k_{\rm B}T_{\rm c}=1$ units). Blue points represent configurations that violate the classical TUR and Pareto fronts to the red dashed line (see legend). Axes are normalized by the respective maximum values $\langle\dot{W}\rangle_{\rm max}=9.5\times 10^{-4}k_{B}T_{\rm c}/\hbar$, ${\rm Var}[\dot{W}]_{\rm max}=5.1\times 10^{-3}(k_{B}T_{\rm c}/\hbar)^{2}$, and $\eta_{C}=0.9$.
Therefore, it becomes clear that the NIC machine is able to overcome the classical equivalent machine in some configurations which are not witnessed by TUR violations as in the case of Hamiltonian-induced coherence between degenerate levels; but not in all configurations. Contrary to the case of Hamiltonian-induced coherence, noise-induced one has not necessarily a beneficial impact in the thermodynamic performance of the NIC machine, whose operation in terms of power, efficiency and reliability can be surpassed in some regimes by a purely classical equivalent machine using the same set of resources.
The region of parameters in which a quantum-thermodynamic advantage over the classical equivalent is verified can be obtained analytically (see App. C) and is given by:
Detrimental effects from coherence are obtained for regions where the above inequality is inverted. In particular, in the symmetric case $\gamma_{\rm c}^{\rm a}=\gamma_{\rm c}^{\rm b}$ and $\gamma_{\rm h}^{\rm a}=\gamma_{\rm h}^{\rm b}$ we will have that $\gamma_{\mathrm{h}}^{\beta}=0$ and the system effectively becomes a three level system without coherence in the steady state. In this limit, the quantum system and its classical analog become indistinguishable. These conclusions are in contrast with some previous claims in the literature claiming quantum enhancements (also in the symmetric case) for same or similar NIC models, see e.g. Refs. [16, 19, 17, 18], which lacked a careful comparison to classical thermodynamic-equivalent models.
VII Pareto optimization and practical advantages
In the previous sections we have seen that only for the case of Hamiltonian-induced coherence, corresponding e.g. to the cases of the three-level amplifier introduced in Sec. II.1 and to the three-qubit refrigerator in Sec. II.2, a quantum-thermodynamic advantage is guaranteed in out-of-equilibrium conditions for all range of parameters. In that case, there is a net reduction of the output current fluctuations as compared to classical equivalent machines operating at same (average) output power and efficiency, hence providing performances not achievable with classical stochastic machines using the same amount of resources. Still in order to proof a practical (and hence useful) advantage, an analysis of the regimes where such an improved reliability is obtained is needed. Here we show that the advantages identified before occur indeed along optimal performance, e.g. for maximum power and for maximum efficiency regimes.
Since optimizing the machine performance involves more than one desired objective (maximize power, maximize efficiency, minimize fluctuations), multi-objective optimization techniques, also called Pareto optimization, are required. These techniques are well known in engineering [93] and have been applied to a wide range of problems, from network theory [34] to betting strategies [35] and phenotypic response [36]. They have also attracted some attention recently in the field of stochastic and quantum thermodynamics, where they have been applied to the optimization of heat engine cycles in the slow-driving regime [37, 38, 39].
In order to solve a multi-objective optimization problem (let’s take for example a two-dimensional problem) we first define a utility function that depends on the quantities to be optimized (e.g. maximization of power and minimization of fluctuations) with the form
$\Omega:=\lambda\langle\dot{W}\rangle-(1-\lambda)\mathrm{Var}[\dot{W}],$ (42)
where the parameter $\lambda$ takes values in the range $[0,1]$. The role of the parameter $\lambda$ is to give weights to the two objectives, thus defining a single optimization problem for the whole function $\Omega$. By maximizing $\Omega$ we obtain solutions that are both optimal in terms of maximizing power and minimizing fluctuations according to a fixed weight $\lambda$. For $\lambda=0$ we recover the problem of only minimizing fluctuations (whose solution is the zero fluctuations point), while in the case $\lambda=1$ we recover the problem of only maximizing power (leading to the maximum power point). Indeed varying $\lambda$ we obtain a family of standard optimization problems for a single variable, $\Omega$. The solution to the multi-objective optimization problem —the so-called Pareto front— is hence constructed by combining the solutions for every single-objective optimization problems, i.e. for all $\lambda\in[0,1]$.
We applied the above optimization technique to the prototypical models introduced in Sec. II for numerically obtaining Pareto fronts to three different multi-objective optimization problems, each of which optimizing two of the three performance indicators of the machine, namely average output power $\langle\dot{W}\rangle$, efficiency $\eta$ and power fluctuations $\mathrm{Var}[\dot{W}]$. For illustrative purposes we focus first on the three-level amplifier of Sec. II.1, and then on the three-qubit autonomous refrigerator of Sec. II.2.
In Fig. 7 we can see the shape of the Pareto front (red curves) of the three-level amplifier for the three different optimization problems with fixed environmental temperatures and varying $\omega_{\mathrm{d}}$, $\gamma_{\mathrm{h}}$, $\gamma_{\mathrm{c}}$ and $g$. There, each circle (independently of its color) is generated from a possible configuration (choice of parameters) of the machine. For the optimization problems involving maximizing power $\langle\dot{W}\rangle$ and minimizing fluctuations $\mathrm{Var}[\dot{W}]$ (Fig. 7a) or maximizing power $\langle\dot{W}\rangle$ and minimizing efficiency $\eta$ (Fig. 7b), the Pareto fronts include a whole branch of solutions indicating a trade-off between the optimization objectives. These go from the zero fluctuations point to the maximum power one in the first case, and from the Carnot point to maximum power point in the second one. On the other hand, for the problem of minimizing fluctuations $\mathrm{Var}[\dot{W}]$ and maximizing efficiency $\eta$ (Fig. 7c) the Pareto front collapses to a single point, or optimal solution, achieving zero fluctuations and Carnot efficiency (but at zero output power).
Figure 8: Fluctuations ratio and two dimensional multi-objective optimization of power $\langle\dot{W}\rangle$, power fluctuations ${\rm Var}[\dot{W}]$ and efficiency $\eta$ for the Coherent three-level amplifier. Points are generated from $10^{5}$ different configurations in the parameter space $\omega_{\mathrm{d}}\in[0.1,4.9]$, $\gamma_{\mathrm{h}/\mathrm{c}}\in[10^{-5},10^{-2}]$ and $g\in[10^{-5},10^{-3}]$ and fixed $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{h}}=0.1$, $\epsilon_{2}=5$ (energies are given in $k_{\rm B}T_{\rm c}=1$ units). The color scale stands for the fluctuation ratio $\mathcal{R}$ and Pareto fronts are highlighted in red (see legend). Axes are normalized by the respective maximum values $\langle\dot{W}\rangle_{\rm max}=9.5\times 10^{-4}k_{B}T_{\rm c}/\hbar$, ${\rm Var}[\dot{W}]_{\rm max}=5.1\times 10^{-3}(k_{B}T_{\rm c}/\hbar)^{2}$, and $\eta_{C}=0.9$.
The role of coherence in the performance of the engine can be unveiled by coloring the circles representing each parameter configuration according to their value of the TUR ratio $\mathcal{Q}_{\mathrm{he}}$ (orange for $\mathcal{Q}_{\mathrm{he}}\geq 2$ and blue for TUR violations, $\mathcal{Q}_{\mathrm{he}}<2$). As can be appreciated in Fig. 7, in all cases the configurations that violate the TUR span the domain classically achievable. More importantly, in Figs. 7a and 7b, we observe TUR violations within the Pareto fronts for large power outputs, and including the maximum power point. This result implies that the coherent three-level amplifier shows a quantum-thermodynamic advantage (as spotted by the violation of the TUR) in relevant regimes of operation, where the machine performs in an optimal way. The TUR violations in Fig. 7 are to be compared with the value of the fluctuations ratio $\mathcal{R}$ for the same parameter configurations in the optimization problems, as shown in Fig. 8 for power and fluctuations (a), and power and efficiency (b). As predicted by Theorem 1 in Sec. V, we obtain positive values $\mathcal{R}>0$ except at zero power, when equilibrium is reached. Moreover, the presence of large values of $\mathcal{R}$ along the Pareto front confirm the presence of practical quantum-thermodynamic advantages in optimal regimes of operation for $\langle\dot{W}\rangle>0$, where the largest values of $\mathcal{R}$ are obtained at the maximum power point.
Figure 9: Fluctuations ratio and two dimensional multi-objective optimization of cooling power $\langle\dot{Q}_{\rm c}\rangle$, its fluctuations ${\rm Var}[\dot{Q}_{\rm c}]$ and the coeficient of performance $\eta$ for the three-qubit autonomous refrigerator. Points are generated from $10^{5}$ different configurations in the parameter space $\epsilon_{1}\in[0.1,4.9]$, $\gamma_{\mathrm{h}/\mathrm{m}/\mathrm{c}}\in[10^{-5},10^{-2}]$ and $g\in[10^{-5},10^{-3}]$ and fixed parameters $\beta_{\mathrm{c}}=1$, $\beta_{\mathrm{m}}/\beta_{\mathrm{c}}=0.5$, $\beta_{\mathrm{h}}/\beta_{\mathrm{c}}=0.05$, $\epsilon_{2}=5$ (energies are given in $k_{\rm B}T_{\rm c}=1$ units). Axes are normalized by the respective maximum values $\langle\dot{Q}_{\rm c}\rangle_{\rm max}=2.9\times 10^{-4}k_{B}T_{\rm c}/\hbar$, ${\rm Var}[\dot{Q}_{\rm c}]_{\rm max}=2.4\times 10^{-3}(k_{B}T_{\rm c}/\hbar)^{2}$, and $\eta_{\rm abs}=0.9$
On the other hand, for the three-qubit autonomous refrigerator, the TUR is never violated and hence it cannot be used to spot practical advantages in optimal operation regimes. In this case, our method is nevertheless able to show that practical quantum-thermodynamic advantages can be obtained despite the absence of TUR violations, as predicted by Theorem 1 in Sec. V. In Fig. 9 we show the colormap of $\mathcal{R}$ for the corresponding two optimization problems of the three-qubit refrigerator with fixed environmental temperatures and varying $\omega_{\mathrm{d}}$, $\gamma_{\mathrm{h}}$, $\gamma_{\mathrm{c}}$ and $g$. The corresponding Pareto fronts are shown again in red color, which have qualitatively similar shapes as in the coherent three-level amplifier [Fig. 8]. Also for this case, Fig. 9 reveals that configurations with values of $\mathcal{R}>0$ are obtained along the Pareto fronts, confirming the presence of quantum-thermodynamic advantages produced by the presence of degenerate coherence in optimal regimes of operation. In this case, maximum values appear along all the Pareto front ${\rm Var[\dot{Q}_{\rm c}]}-\langle Q_{\rm c}\rangle$. However, we observe that the values of $\mathcal{R}$ are in general much smaller than in the previous case, showing a higher impact of energetic coherence to produce quantum advantages in comparison to degenerate one.
VIII Conclusions
We have characterized the thermodynamic impact of Hamiltonian and noise-induced coherence in the performance of quantum thermal machines operating in nonequilibrium steady states. While machines displaying Hamiltonian coherence in their steady states lead to quantum-thermodynamic advantages in terms of their tradeoff between power, efficiency and stability, this is not always the case for machines displaying noise-induced coherence. These results, obtained through direct comparison of generic quantum thermal machines with axiomatically-constructed classical equivalent models, ensure the presence of quantum-enhanced performance for models of Hamiltonian-induced coherence in out of equilibrium situations. On the other hand, they also imply that several previous claims of thermodynamic advantages in machines with noise-induced coherence may be spurious, and suggests that they should be reassessed on a case-by- case basis, by determining the set of parameters in which quantum advantages may really exist (if any). Indeed, even if the dynamics of these machines may exhibit purely quantum features, a classical Markovian machine using the same set of resources (energy gaps and bath temperatures) can be systematically constructed that performs just as well or better than the original machine.
The classical equivalent model employed here allowed us a comparison of the current fluctuations between a quantum device and a classical (incoherent) counterpart that outputs the same average currents while employing identical thermodynamic resources. This notion of classical equivalent machine constructs on the idea of emulability discussed in Ref. [23], and is accompanied by a general methodology for its derivation in generic cases. The classical equivalent can be constructed for virtually any quantum steady-state machine working in the weak-coupling regime, and under weak-driving conditions, namely, when the driving can be considered a perturbation of the (bare) machine Hamiltonian. Extensions of this method to the case of strong couplings or strong periodic driving (e.g. using Floquet formalism) are an interesting direction for future research which may allow addressing quantum-thermodynamic advantages in a larger class of quantum devices.
Using the three-level amplifier and the three-qubit-fridge as main examples of a quantum thermal machines displaying Hamiltonian coherence (the first energetic coherence and the second degenerate one), we have illustrated situations where coherence is always beneficial. In these cases the performance match or exceeds the performance of its classical equivalent counterpart in all possible configurations, with improvements in the machine stability that become greater far from equilibrium. In the case of the three-level amplifier, the parameter regions where these improvements are maximal coincide indeed with regimes where the thermal machine breaks the TUR bound. However, we also observe wide regions showing (smaller) thermodynamic improvements that are not witnessed by TUR violations. This is indeed entirely the case for the three-qubit-fridge, where the TUR by itself is not able to spot thermodynamic enhancements. Importantly, our results reveal that in order to observe a genuine quantum-thermodynamic advantage, it is necessary (and sufficient) to consider the fluctuations in the currents (at least at the level of the variance), in contrast to previous assessments based on more limited notions of classical equivalents (leading also to a less stringent comparison) [9].
Beyond the direct comparison between quantum machines and their respective equivalents for given parameters, we also provided an all-to-all comparison computing the optimal configurations (Pareto front) of quantum and classical machines maximizing power and efficiency while minimizing fluctuations. Our analysis reveals that quantum-thermodynamic advantages in both the three-level amplifier and the three-qubit-fridge occur indeed within the optimal front, allowing thus for a practical quantum-thermodynamic advantage. However, it is also clear from our analysis that in the three-qubit-fridge, the advantages are much lower than in the three-level amplifier. Since implementations of these or similar thermal machine models have been proposed in a number of platforms [58, 87, 57, 13, 56], and have been actually experimentally realized with NV-centers in diamond [12] and trapped ions [11] respectively, it would be interesting to test our results for the quantum-thermodynamic advantages reported here in relevant regimes by suitably measuring the variance of the currents [94, 95].
Acknowledgements.
We thank Roberta Zambrini for comments and interesting discussions. We wish to acknowledge support from the María de Maeztu project CEX2021-001164-M for Units of Excellence and CoQuSy project PID2022-140506NB-C21, funded by MICIU/AEI/10.13039/501100011033/FEDER, UE. GM acknowledges the Ramón y Cajal program RYC2021-031121-I funded by MICIU/AEI/10.13039/501100011033 and European Union NextGenerationEU/PRTR. JAAM acknowledges the Conselleria d’Educació, Universitat i Recerca of the Balearic Islands (Grant FPI FPI_$\_058\_2022$).
Appendix A General expression of current first moments with local dissipation
In this appendix we derive a general expression for the average heat currents including the cases in which coherence may be generated either by Hamiltonian or noise-induced sources. We start by expanding the expression for th standard definition for the heat currents as given in Eq. (13), that is:
$\begin{split}&\langle\dot{Q}_{r}\rangle=\sum_{k}\operatorname{Tr}[(H_{0}+V(t))\mathcal{D}_{k}^{(r)}[\pi(t)]]\\ &=\sum_{k}\operatorname{Tr}[H_{0}\mathcal{D}_{k}^{(r)}[\pi(t)]]=\sum_{k}\operatorname{Tr}[\mathcal{D}_{k}^{\dagger(r)}[H_{0}]\pi(t)]\\ &=\sum_{k}\sum_{m,l}\bra{m}L_{k}^{(r)\dagger}H_{0}L_{k}^{(r)}-\dfrac{1}{2}\{L_{k}^{(r)\dagger}L_{k}^{(r)},H_{0}\}\ket{l}\pi_{lm}\end{split}$ (43)
where the terms $\operatorname{Tr}[V(t)\mathcal{D}_{k}^{(r)}[\pi(t)]]\sim g\gamma_{0}^{k}$ have been neglected in accordance to the weak coupling and weak driving approximations. We have applied the cyclic property of the trace to obtain the adjoint dissipator $\mathcal{D}_{k}^{\dagger(r)}[\cdot]:=L_{k}^{(r)\dagger}\cdot L_{k}^{(r)}-\dfrac{1}{2}\{L_{k}^{(r)\dagger}L_{k}^{(r)},\cdot\}$. Now, using the explicit form of the Lindblad operators in Eq. (4) and the bare Hamiltonian, $H_{0}=\sum_{i}\epsilon_{i}\ket{i}\bra{i}$ |, we can obtain each term in the sum of the last expression:
@@ Appendix A General expression of current first moments with local dissipation
where the $\delta$ functions arise from the fact that the Lindblad operators produce jumps only between levels with a fixed energy gap $\Delta\epsilon_{k}=\pm\Delta\epsilon_{r}$, determined by the reservoirs and we used $\delta(\Delta\epsilon_{ji}-\Delta\epsilon_{k})\delta(\Delta\epsilon_{jn}-\Delta\epsilon_{k})=\delta(\Delta\epsilon_{ji}-\Delta\epsilon_{k})\delta(\Delta\epsilon_{ji}-\Delta\epsilon_{ij})=\delta(\Delta\epsilon_{ji}-\Delta\epsilon_{k})\delta(\Delta\epsilon_{ni})$. By introducing the expressions in (44) into (43) we arrive at a general expression for the heat currents valid for both degenerate and non-degenerate level structures in the machine:
$\begin{split}\langle\dot{Q}_{r}\rangle=\sum_{k,n,i,j}\sqrt{\gamma_{ij}\gamma_{nj}}\delta(\Delta\epsilon_{ji}-\Delta\epsilon_{k})\delta(\Delta\epsilon_{ni})\\ \times[\epsilon_{j}\pi_{ni}-\epsilon_{n}\mathrm{Re}(\pi_{ni})].\end{split}$ (45)
$\begin{split}\langle\dot{Q}_{r}\rangle=&\sum_{k,i,j}\delta(\Delta\epsilon_{ji}-\Delta\epsilon_{k})(\epsilon_{j}-\epsilon_{i})\gamma_{ij}\pi_{ii}\\ &=\sum_{i<j}(\epsilon_{j}-\epsilon_{i})(\gamma_{ij}\pi_{ii}-\gamma_{ji}\pi_{jj}),\\ \end{split}$ (46)
as given in Eq. (22). On the other hand, for the cases with degenerate energy levels, $\delta(\epsilon_{ni})$ can be zero even for $n\neq i$, and extra terms are obtained. In the case of a single pair of degenerate levels |$\ket{\mathrm{u}}$ and |$\ket{\mathrm{v}}$ we obtain:
$\begin{split}\langle\dot{Q}_{r}\rangle=&\sum_{i<j}(\epsilon_{j}-\epsilon_{i})(\gamma_{ij}\pi_{ii}-\gamma_{ji}\pi_{jj})\\ &+2\sum_{j}(\epsilon_{j}-\epsilon_{\mathrm{v}})(\gamma_{\mathrm{u}j}+\gamma_{\mathrm{v}j})\mathrm{Re}(\pi_{\mathrm{uv}})\end{split}$ (47)
Appendix B Full Counting Statistics

### v4

@@ Appendix D Details on the construction of classical equivalent models
We obtain the equations of motion for the density operator elements from the above master equation in Lindblad form, from which we can obtain the expression for a generic element of the density operator:
$\begin{split}\dot{\rho}_{ij}=-\mathrm{i}\bra{i}[V_{I},\rho]\ket{j}+\sum_{r}\sum_{k}\bra{i}\mathcal{D}_{k}^{(r)}[\rho]\ket{j},\end{split}$ (50)
where for the ease of simplicity in the notation, we ignore the prime symbols on the density matrix elements. After expanding the form of the dissipators $\mathcal{D}_{k}^{(r)}[\rho]=L_{k}^{(r)}\rho L_{k}^{(r)\dagger}-\frac{1}{2}\{L_{k}^{(r)\dagger}L_{k}^{(r)},\rho\}$, we obtain the following terms contributing to the above expression:
$\bra{i}[V_{I},\rho]\ket{j}=$ $\Delta_{\rm d}(\delta_{j\rm u}\rho_{i\rm u}+\delta_{i\rm u}\rho_{\rm uj})+g(\delta_{i\mathrm{u}}\rho_{\mathrm{v}j}$
$+\delta_{i\mathrm{v}}\rho_{\mathrm{u}j}-\delta_{\mathrm{v}j}\rho_{i\mathrm{u}}-\delta_{\mathrm{u}j}\rho_{i\mathrm{v}}),$
$\bra{i}L_{k}^{(r)}\rho L_{k}^{(r)\dagger}\ket{j}=$ $\sum_{n,m}\alpha_{ni}^{k}\alpha_{mj}^{k}\sqrt{\gamma_{ni}\gamma_{mj}}\rho_{nm},$
$\bra{i}L_{k}^{(r)\dagger}L_{k}^{(r)}\rho\ket{j}=$ $\sum_{n,m}\alpha_{im}^{k}\alpha_{nm}^{k}\sqrt{\gamma_{im}\gamma_{nm}}\rho_{nj},$ (51)
$\bra{i}\rho L_{k}^{(r)\dagger}L_{k}^{(r)}\ket{j}=$ $\sum_{n,m}\alpha_{nm}^{k}\alpha_{jm}^{k}\sqrt{\gamma_{nm}\gamma_{jm}}\rho_{in},$
@@
where we can again use the fact that the Lindblad operators produce jumps only between levels with a fixed energy gap $\Delta\epsilon_{k}=\pm\Delta\epsilon_{r}$ and rewrite the $\alpha$-terms as delta functions:
In the case of Hamiltonian-induced coherence the $\delta$ functions in (52) can be simplified by taking into account the fact that in this case different transitions cannot have the same $\Delta\epsilon_{k}$ associated with them. Then (52) becomes:
$\begin{split}&\alpha_{ni}^{k}\alpha_{mj}^{k}=\delta(\Delta\epsilon_{in}-\Delta\epsilon_{k})\delta_{ij}\delta_{nm},\\ &\alpha_{im}^{k}\alpha_{nm}^{k}=\delta(\Delta\epsilon_{mi}-\Delta\epsilon_{k})\delta_{ni},\\ &\alpha_{nm}^{k}\alpha_{jm}^{k}=\delta(\Delta\epsilon_{mn}-\Delta\epsilon_{k})\delta_{nj},\end{split}$ (53)
which results in the set of equations (III.1).
In the case of noise-induced coherence, we don’t have an interaction Hamiltonian, and therefore $V_{I}=0$ in the first expression on (D), which doesn’t need to be considered. The $\delta$ functions in (52), on the other hand, have more terms, since there are now different transitions associated with the same $\Delta\epsilon_{k}$. Now (52) reads:
$\alpha_{ni}^{k}\alpha_{mj}^{k}=$ $\delta(\Delta\epsilon_{in}-\Delta\epsilon_{k})(\delta_{ij}\delta_{nm}+\delta_{i\mathrm{u}}\delta_{j\mathrm{v}}\delta_{nm}+\delta_{i\mathrm{v}}\delta_{j\mathrm{u}}\delta_{nm}$
$+\delta_{n\mathrm{u}}\delta_{m\mathrm{v}}\delta_{ij}+\delta_{n\mathrm{v}}\delta_{m\mathrm{u}}\delta_{ij}),$ (54)
$\alpha_{im}^{k}\alpha_{nm}^{k}=$ $\delta(\Delta\epsilon_{mi}-\Delta\epsilon_{k})(\delta_{ni}+\delta_{i\mathrm{v}}\delta_{n\mathrm{u}}+\delta_{i\mathrm{u}}\delta_{n\mathrm{v}}),$
$\alpha_{nm}^{k}\alpha_{jm}^{k}=$ $\delta(\Delta\epsilon_{mn}-\Delta\epsilon_{k})(\delta_{nj}+\delta_{j\mathrm{v}}\delta_{n\mathrm{u}}+\delta_{j\mathrm{u}}\delta_{n\mathrm{v}}).$
In view of (D) the equations of motion are:
for the levels $n\neq\{\text{u},\text{v}\}$, and:
$\begin{split}\dot{\rho}_{nn}=&\sum_{i}\left[\gamma_{in}\rho_{ii}-\left(\gamma_{ni}\rho_{nn}+\sqrt{\gamma_{\text{u}i}\gamma_{\text{v}i}}\text{Re}(\rho_{\text{uv}})\right)\right],\\ \dot{\rho}_{\text{uv}}=&\sum_{i}[2\sqrt{\gamma_{i\text{u}}\gamma_{i\text{v}}}\rho_{ii}-\sqrt{\gamma_{\text{u}i}\gamma_{\text{v}i}}(\rho_{\text{uu}}+\rho_{\text{vv}})\\ &-(\gamma_{\text{u}i}+\gamma_{\text{v}i})\rho_{\text{uv}}],\\ \end{split}$ (56)
Finally, it can be shown that all energy currents in the quantum machine and the classical equivalent are equal, also in the case of noise-induced coherence. This is a consequence of the fact that in the steady state, the value of the coherence can be written entirely in terms of the steady-state populations, as given in Eq.  (12). That leads to the substitution of the original collective transitions by independent ones with modified rates, Eq.  (13). As a consequence, the expression for the original heat currents in the collective transitions, Eq.  (11), can be also expressed as a function of the new transitions using the same steady-state populations, therefore guaranteeing the same average currents in the original and classical thermodynamic equivalent models.
Appendix E Full Counting Statistics methods